Unpolarized light of intensity Io is incident on the following arrangements of polarizing sheets. The first polarizer is oriented vertically; the second is oriented so it makes an angle of 30o with the vertical; the third and last polarizer makes an angle of -30o with the vertical. Determine the intensity of the light transmitted through this group of polarizers as a fraction of Io.


1
Expert's answer
2021-05-19T10:47:45-0400

According to the Malus's law, the intensity of the light after it passes the second polarizer is the following:


"I_1 = I_0\cos^2\theta_1"

where "\theta_1 = 30\degree" is the agle between the first and the second polarizer.

Similarly, the intensity of the light after it passes the third polarizer is the following:


"I_2 = I_1\cos^2\theta_2"

where "\theta_2 = 60\degree" is the agle between the second and the third polarizer.

Finally, obtain:


"I_2 = I_0\cos^2\theta_1\cos^2\theta_2\\
\dfrac{I_2}{I_0} = \cos^2\theta_1\cos^2\theta_2\\
 \dfrac{I_2}{I_0} = \cos^230\degree\cdot \cos^260\degree =\left(\dfrac{\sqrt{3}}{2}\right)^2\cdot\left(\dfrac{1}{2} \right)^2 = 0.1875"

Answer. The intensity of the light transmitted through this group of polarizers as a fraction of Io is 0.1875.